OpenStudy (anonymous):

Find the coordinates of the foci for the ellipse: (x + 3)2/9 + (y - 1)2/25 = 1 A. (3, -1) and (3, 5) B. (-3, 3) and (-3, 1) C. (-3, 5) and (-3, -3) D. (1,1) and (-7,1)

OpenStudy (anonymous):

http://www.mathopenref.com/ellipsefoci.html Check this out...should help you understand how to find it.

OpenStudy (anonymous):

@Chris9k Nope still not able to figure out :/

OpenStudy (anonymous):

You have your center at (-3, 1) that it is in standard form and you can pull that straight out of the formula of an ellipse. Do you see where that came from?

OpenStudy (anonymous):

@Chris9k Yes I think so, did you get the -3 from multiplying the 3 and -1

OpenStudy (anonymous):

You know y is your major axis because 25>9....a^2 =25. b^2 =9. Then you take the square root of your major axis squared minus your minor axis squared. That give you sqrt(5^2 - 3^2) = sqrt(25-9) = sqrt(16) = 4. The foci will be along the major axis which is the y-axis. You then + and - 4 from the y value of the center. (-3, 1+4) and (-3,1-4) = (-3,5) and (-3, -3).

OpenStudy (anonymous):

I got the center from (x+3) part of the equation....that means the x-term of your center is -3. for the y-term of the center you go from (y-1) which means the y value of the center is +1...so (-3,1) for center.

OpenStudy (anonymous):

@Chris9k Ooh Ok I got it, Thank You very much

OpenStudy (anonymous):

Great! You're welcome.